OpenStudy (anonymous):

which of the following functions is represented by the Maclaurin series sum from n=0 to infinity [(-9)^n x^(4n)]/(2n)! a) cos(3x^2) b)sin(3x^2) c)cos(9x^4) d)arctan(3x) e)e^(-3x^2)

myininaya (myininaya):

So we know that the Maclaurin series is a taylor series expansion of a function about 0. That is \[f(x)=f(0)+f'(0)x+\frac{f''(0)}{2!}x^2+\frac{f^{(3)}(0)}{3!}x^3+...+ \frac{f^{(n)}(0)}{n!}x^n+...\]

myininaya (myininaya):

We are given the expansion and we are asked to find the function The expansion is: \[1+\frac{-9}{2}x^4 +\frac{81}{4!}x^{8}+\frac{-729}{6!}x^{12}+...\]

myininaya (myininaya):

You could eliminate some of these by evaluating f(0) If you don't get 1 then throw it out

myininaya (myininaya):

Are you there?

OpenStudy (anonymous):

yes sorry!

myininaya (myininaya):

Which choices can we go ahead and eliminate?

myininaya (myininaya):

We are given f(0)=1 which of the functions given satisfy that.

OpenStudy (anonymous):

a c and e

OpenStudy (anonymous):

so we can eliminate b and d

myininaya (myininaya):

yep

myininaya (myininaya):

now looking at a can you find it's fifth term which is \[\frac{f^{(4)}(0)}{4!}x^4\]

myininaya (myininaya):

we are trying to see if this is the same as \[\frac{-9}{2}x^4 \]

myininaya (myininaya):

you are going to need the 4th derivative

OpenStudy (anonymous):

okay ill calculate that

OpenStudy (anonymous):

I dont think that a will be the answer because these derivatives are getting super long

OpenStudy (anonymous):

i got the fourth derivative but if i plug in 0 i get 0

OpenStudy (anonymous):

i did that wrong, now i got -108 after plugging in

myininaya (myininaya):

\[f(x)=\cos(3x^2)\] \[f'(x)=-6x \sin(3x^2) \] \[f''(x)=-6\sin(3x^2)-6x(6x)\cos(3x^2)=-6\sin(3x^2)-36x^2\cos(3x^2)\] So fall all of these will give us 0 when pluggin in 0 which is a good sign since in the other series we don't see another term until x^4 \[f^{(3)}(x)=-36xcos(3x^2)-72xcos(3x^2)+36(6x)\sin(3x^2) \\ =-36xcos(3x^2)-72xcos(3x^2)+216xsin(3x^2)\] still have zero when we plug in 0 ok...I will stop writting what I was writing yep f^(4)(0)=-108

myininaya (myininaya):

so that means the fourth term is?

myininaya (myininaya):

remember we are trying to compare it to \[\frac{-9}{2}x^4 \]

OpenStudy (anonymous):

-108/4! X^4

myininaya (myininaya):

and does -108/4! simplify

myininaya (myininaya):

or reduce

OpenStudy (anonymous):

yes it simplifies to -9/2

OpenStudy (anonymous):

so a is the answer!

myininaya (myininaya):

It looks that way. I would look at the other 4terms from the other functions but that seems to be a long process

OpenStudy (anonymous):

thanks for your help!! I really apprectiate it!

myininaya (myininaya):

np